### v1

# Superposition in Graph Neural Networks

Abstract
Interpreting graph neural networks (GNNs) is difficult because message passing mixes signals and internal channels rarely align with human concepts. We study superposition, the sharing of directions by multiple features, directly in the latent space of GNNs. Using controlled experiments with unambiguous graph concepts, we extract features as (i) class-conditional centroids at the graph level and (ii) linear-probe directions at the node level, and then analyze their geometry with simple basis-invariant diagnostics. Across GCN/GIN/GAT we find: increasing width produces a phase pattern in overlap; topology imprints overlap onto node-level features that pooling partially remixes into task-aligned graph axes; sharper pooling increases axis alignment and reduces channel sharing; and shallow models can settle into metastable low-rank embeddings. These results connect representational geometry with concrete design choices (width, pooling, and final-layer activations) and suggest practical approaches for more interpretable GNNs.
1 Introduction
Understanding what features a model represents and how they are arranged in latent space is central to trustworthy ML. For graph neural networks (GNNs), this is unusually hard. Unlike pixels or tokens, graphs do not offer a fixed coordinate system across inputs; receptive fields are relational and variable; and many signals are structural (motifs, roles, spectral patterns) rather than obviously human readable. As a result, most GNN interpretability work answer which nodes or edges mattered, not what internal features the network formed or how those features are arranged (Ying et al., 2019; Luo et al., 2020; Yuan et al., 2022; Kakkad et al., 2023).
A key obstacle to interpretability is superposition: many features are packed in fewer directions, creating polysemantic channels and entangled geometries (Elhage et al., 2022; Scherlis et al., 2023). Superposition has been studied in MLPs and transformers, but its behavior in GNNs, where message passing and pooling constrain geometry, remains underexplored. This motivates the central theme of our work, which asks: How does superposition arise in GNNs, how do architectural and graph-structural choices modulate it, and what are the downstream consequences for interpretability?
To answer these questions, we build small, controllable datasets where the relevant graph concepts are unambiguous (e.g., "two adjacent identical types" or "contains a triangle and a hexagon"), train standard GNNs, and look directly at the geometry of their internal representations – both at the node level (before pooling) and at the graph level (after pooling). We then use two simple, basis-invariant diagnostics: one that says "how many distinct axes are effectively used" and another that says "how tightly packed the directions are compared with an ideal arrangement."
Our study yields several main findings:
  • Width produces a phase pattern. As the final hidden layer widens, feature overlap first decreases, then briefly increases around the point where capacity matches the number of concepts, then decreases again. At high width, graph-level features approach near-ideal packing, but node-level concepts often remain entangled – evidence that readout layers can disentangle what the message-passing stack keeps mixed.
  • Pooling sharpness encourages axis alignment. Making pooling more "winner-take-all" (e.g., max pooling) increases alignment to coordinate axes and reduces node-level feature sharing. We explain this with two complementary arguments: gradients concentrate on large coordinates, and under noise the axis-aligned choice loses less information than an oblique one.
  • Rank collapse can appear even in shallow GNNs. In some runs the numerical rank of the pooled representation stays below the number of learned features while the accuracy remains high. We link this to hard gating after aggregation (exactly zeroing channels) and to a loss-driven preference for "mutually obtuse" class directions that resists activating new dimensions.
Contributions. (i) A representation-centric framework for analyzing superposition in GNNs that works across architectures and layers; (ii) robust, simple diagnostics and feature extraction procedures (centroids and probes) that avoid assumptions about neurons or coordinates; (iii) empirical maps of how width, topology, and pooling shape superposition; and (iv) practical guidance – when to expect entanglement, when pooling helps, and how to avoid low-rank traps.
2 Background
2.1 Superposition and the interpretability of internal representations
Classic unit analyses in CNNs found channels that act like concept detectors (Zeiler and Fergus, 2014; Zhou et al., 2014; Bau et al., 2017; Olah et al., 2017); related work spans RNNs and transformers (Karpathy et al., 2015; Strobelt et al., 2017; Gurnee et al., 2023, 2024; Geva et al., 2020), and early GNN studies exist (Xuanyuan et al., 2023). However, concepts need not be axis-aligned: they can correspond to arbitrary directions or nonlinear detectors. Directional tools (TCAV, linear probes) therefore treat a feature as a vector in latent space (Kim et al., 2018; Alain and Bengio, 2016); recent sparse-autoencoder work aims to demix polysemantic units (Bricken et al., 2023; Gao et al., 2024).
Superposition is the phenomenon where models represent more features than available dimensions by allowing features to share neurons/directions, producing polysemantic units and entangled directions (Elhage et al., 2022). Both neuron-level and mechanistic analyses benefit from (approximate) decomposability: having independently meaningful, linearly separable features. Superposition violates this assumption, obscuring neuron-level semantics and complicating circuit reconstruction. Methods that demix features (e.g., sparse autoencoders) can mitigate these issues (Bricken et al., 2023; Gao et al., 2024).
2.2 Representations in GNNs
A message-passing layer aggregates neighbor information and applies a local transform,
$\mathbf{H}^{(l+1)}=\phi\!\left(\textsc{Agg}(\mathbf{H}^{(l)},\mathbf{A}),\,\mathbf{W}^{(l)}\right),$ (1)
with architecture-specific aggregation (e.g., normalized sum for GCN (Kipf and Welling, 2017), unnormalized sum + MLP for GIN (Xu et al., 2019), attention for GAT Veličković et al. (2018)). After $L$ layers, node embeddings $\mathbf{H}^{(L)}$ are pooled into a graph embedding $\mathbf{h}_{G}=g(\mathbf{H}^{(L)})$ via sum/mean/max.
Superposition pressure in GNNs. Message passing repeatedly mixes many neighbor features into fixed-width hidden states. This creates capacity pressure that encourages superposition – multiple features sharing the same channels – leading to polysemantic units and entangled directions. While superposition has been analyzed in toy feedforward/transformer settings, to our knowledge there is little analogous work for GNNs. This gap further complicates neuron-level semantics and the reconstruction of GNN circuits.
3 Quantifying Superposition in Representations
3.1 Interpreting and extracting features
Prior work often reasons about superposition at the input to layers (e.g., pixels/tokens). For GNNs, we study it where it matters for interpretability: in the latent spaces of nodes and graphs. We ask: how many features are packed into how few directions, and how well are those directions separated?
Following (Alain and Bengio, 2016; Kim et al., 2018; Elhage et al., 2022; Bricken et al., 2023; Gao et al., 2024), we treat a feature as a direction in a model’s latent space that supports a prediction. For a GNN with forward pass
$\mathbf{X}\xrightarrow{\phi_{1}}\mathbf{H}^{(1)}\xrightarrow{\phi_{2}}\cdots\xrightarrow{\phi_{L}}\mathbf{H}^{(L)}\rightarrow\mathbf{h}_{G}\xrightarrow{g}\hat{\mathbf{y}}$,
each $\phi_{\ell}$ is a message-passing block, $\mathbf{H}^{(\ell)}\!\in\!\mathbb{R}^{n_{\ell}\times d_{\ell}}$ are node embeddings, $g$ is a graph-pool, and $\mathbf{h}_{G}\!\in\!\mathbb{R}^{d}$ is the graph embedding. We interpret features in two complementary ways.
(1) Linear-probe features (model-decodable concepts).
Given embeddings $\{\mathbf{h}_{G}\}$ (graph-level) or rows of $\mathbf{H}^{(\ell)}$ (node-level) and binary targets $y_{\ell}\!\in\!\{0,1\}$ for concept $\ell$, we fit a logistic probe on a held-out split:
$s_{\ell}=\mathbf{w}_{\ell}^{\top}\mathbf{z}+b_{\ell},\quad\mathbf{z}\in\{\mathbf{h}_{G}\}\ \text{or}\ \{\mathbf{H}^{(\ell)}_{i:}\}$.
The probe normal $\mathbf{w}_{\ell}$ (unit-normalized) is taken as the feature direction for concept $\ell$. Since all our geometry is directional, the intercept $b_{\ell}$ is irrelevant.
(2) Class-conditional centroids (task-aligned concepts). On a held-out split, we form ground-truth one-hot sets
$\mathcal{S}_{\ell}^{\mathrm{GT}}=\{G:\mathbf{y}(G)=\mathbf{e}_{\ell}\},\qquad\mathbf{c}_{\ell}=\frac{1}{|\mathcal{S}_{\ell}^{\mathrm{GT}}|}\sum_{G\in\mathcal{S}_{\ell}^{\mathrm{GT}}}\mathbf{h}_{G},$
and stack rows to obtain $C\in\mathbb{R}^{K\times d}$. In the linear/LDA regime, discriminants align with class means (Alain and Bengio, 2016), so centroid directions approximate task-discriminative axes and aggregate the combined effects of message passing and pooling.
Active features. Not all labels are reliably learned in every seed, thus including non-existent or highly confused features would distort superposition measurements. We therefore define an active feature $\ell$ using only the held-out split:
  • Probes: a concept $\ell$ is active if $\mathrm{AUC}_{\ell}\geq 0.60$.
  • Centroids: class $\ell$ is active if in-class recall $\geq 0.5$ and all off-diagonal recalls <0.5<0.5.
Let $\mathcal{L}_{\mathrm{act}}$ be the active set ($|\mathcal{L}_{\mathrm{act}}|=k_{a}$). We form the feature matrix by stacking unit vectors: either $\{\hat{\mathbf{w}}_{\ell}\}_{\ell\in\mathcal{L}_{\mathrm{act}}}$ (probes) or $\{\hat{\mathbf{c}}_{\ell}\}_{\ell\in\mathcal{L}_{\mathrm{act}}}$ (centroids). In the main results we use class-conditional centroids plus linear probes as complementary views: centroids assess task-aligned geometry, probes assess model-decodable directions.
3.2 Proposed Metrics
We propose metrics to measure superposition for a $k_{a}\times d$ feature matrix $C$.
Effective rank (EffRank).
With singular values $\sigma_{1}\geq\sigma_{2}\geq\cdots$ and $p_{i}=\sigma_{i}/\sum_{j}\sigma_{j}$, the entropy-based effective rank is
$\mathrm{EffRank}(C)=\exp\!\Bigl{(}-\sum_{i}p_{i}\log p_{i}\Bigr{)}.$ (2)
It estimates the number of effectively used axes. For centroids we COM-center $C^{\circ}=C-\mathbf{1}\bar{\mathbf{c}}^{\top}$ with $\bar{\mathbf{c}}=\tfrac{1}{k_{a}}\sum_{\ell}\mathbf{c}_{\ell}$ before computing EffRank, removing a global bias direction that no practitioner would interpret as a mechanism. For probe normals we use the raw $C$.
Superposition Index (SI).
To express features per effective axis we use the absolute index
$\mathrm{SI}=\frac{k_{a}}{\mathrm{EffRank}(C)}.$ (3)
$\mathrm{SI}=1$ indicates no extra sharing beyond having $k_{a}$ independent directions; $\mathrm{SI}>1$ means multiple features share axes (greater superposition). This avoids dependence on the ambient $d$.
Welch-Normalized Overlap (WNO). While SI measures the mean number of features distributed over each axis, it does not consider their angular geometry [^1] For example, two different arrangements of packing 6 vectors into 2D: (i) arranged as a near-regular hexagon, and (ii) arranged such that there are two antipodal clusters. Both have $\text{EffRank}\approx 2$.. Therefore to complement SI, we introduce an angular overlap metric called the Welch-normalized overlap: let $\widetilde{C}$ be a row-unit–normalized projection of $C$ and define
$\overline{\cos^{2}}=\frac{2}{k_{a}(k_{a}-1)}\sum_{i<j}\langle\tilde{\mathbf{c}}_{i},\tilde{\mathbf{c}}_{j}\rangle^{2}.$
We compare $\overline{\cos^{2}}$ to (i) the random baseline $1/d_{\mathrm{eff}}$ (independent unit vectors in $\mathbb{R}^{d_{\mathrm{eff}}}$) and (ii) the Welch lower bound $\mu_{*}^{2}=\max\bigl{(}0,\tfrac{k_{a}-d_{\mathrm{eff}}}{d_{\mathrm{eff}}(k_{a}-1)}\bigr{)}$, and report
$\mathrm{WNO}=1-\frac{\tfrac{1}{d_{\mathrm{eff}}}-\overline{\cos^{2}}}{\tfrac{1}{d_{\mathrm{eff}}}-\mu_{*}^{2}}\,$, (4)
so $\mathrm{WNO}=0$ is Welch-optimal (least overlap), $\mathrm{WNO}=1$ is random, and $\mathrm{WNO}>1$ is worse-than-random packing. Since ambient $d$ can be much larger than the subspace actually used. We therefore report intrinsic WNO: set $r=\lceil\mathrm{EffRank}(C^{\dagger})\rceil$ (with $C^{\dagger}=C^{\circ}$ for centroids and $C^{\dagger}=C$ for probes), project $C$ to the top-$r$ right-singular subspace to obtain $C_{r}$, row-normalize, and take $d_{\mathrm{eff}}=r$ in (4). For centroids and graph-level probes we additionally remove a single shared offset direction before the SVD (PC1 removal); for node-level probes we do not remove PC1.222PC1 removal is appropriate when a single global bias dominates all classes (e.g., pooled magnitude); removing it improves sensitivity to relative packing. For node probes this shared offset is not guaranteed and we keep the raw geometry. If $r\leq 1$ or $k_{a}\leq 1$, WNO is undefined and we report NA.
Invariances and scope.
EffRank and WNO are invariant to right-multiplication by orthogonal matrices (basis changes) and to uniform rescaling of rows; they probe geometry intrinsic to the representation rather than coordinate choices. SI quantifies features-per-axis pressure; WNO quantifies how tightly those axes are packed relative to principled baselines.
4 Effects of Model Width and Graph Topology
4.1 Datasets
The Pairwise dataset. We introduce a dataset where graphs are equal length linear chains. Each node has an $n$-dimensional one-hot/zero feature. Class $k$ is active if two adjacent nodes both activate dimension $k$, yielding sparse multi-class targets. The task is attribute driven: since each class is associated with identical graph structure, we can probe for superposition whilst factoring out the effects of varying topology. Sparsity is controlled by the activation probability $p$ of a node having a one-hot (non-zero) feature.
The Conjunction dataset. To isolate topology–driven superposition from effects of extraneous input attributes, we introduce a graph family in which all node features are initialized to the node’s degree, and supervision otherwise depends only on the presence of simple cycles. Let $C_{\ell}(G)$ denote the indicator that graph $G$ contains at least one simple cycle of length $\ell$. We instantiate four motif detectors $\ell\in\{3,4,5,6\}$ (triangles, squares, pentagons, hexagons) and define two labels by conjunctions of cycles:
$y_{A}(G)\;=\;\mathbf{1}\!\{\,C_{3}(G)\wedge C_{6}(G)\,\},\qquad y_{B}(G)\;=\;\mathbf{1}\!\{\,C_{4}(G)\wedge C_{5}(G)\,\}.$
All other graphs receive $y_{A}=y_{B}=0$. Thus the same lower-level motif concepts $\{C_{3},C_{4},C_{5},C_{6}\}$ are re-used across the two tasks, but each label requires a distinct pair. This cleanly separates (i) the formation of topological features from (ii) how a model combines them. A graph is formed by sampling required counts for the four motifs according to the label pattern and then optionally adding extra copies of each motif with independent probability $p_{\text{extra}}$. Motif instances are placed on disjoint node sets and connected by random bridges to avoid trivial overlaps. Because there are no informative node attributes, any successful solution must construct the motif detectors purely via message passing.
4.2 Width-induced superposition
We study how width $d$ affects superposition on PAIRWISE ($k=16$). The first hidden layer is set to the input width ($k$) and the second layer to $d$, inducing a bottleneck when $d<k$. We measure (a) class-conditional centroids in the graph space and (b) node-level concepts $\mathsf{Is}_{t}$ and $\mathsf{NextTo}_{t}$ via linear probes on the last pre-pooling layer: let $x_{i}\in\{e_{1},\ldots,e_{k}\}$ be the one-hot node type and $\mathcal{N}_{i}$ the neighbors of $i$. For type $t$,
$\mathsf{Is}_{t}(i):=\{x_{i}=e_{t}\},\qquad\mathsf{NextTo}_{t}(i):=\mathbf{1}\{\exists j\in\mathcal{N}_{i}:\,x_{j}=e_{t}\}.$
We keep only active features: centroids with in-class recall >0.5>0.5 and off-class recall <0.5<0.5, and probes with AUROC >0.6>0.6. We report intrinsic WNO (computed in the EffRank-selected subspace). Centroids are COM-centered for EffRank; probe normals are not.
(a) Class-conditional centroids of the graph embeddings.
(b) Linear probe features of the $\mathsf{NextTo}$ concept family after the second message-passing layer.
(c) Linear probe features of the $\mathsf{Is}$ concept family after the second message-passing layer.
Figure 1: Superposition effects across bottleneck dimensions $d$. A dashed line corresponds to $k=d$.
[…]
Architecture trends follow inductive bias. For GIN, the post-aggregation MLP tends to funnel many labels into shared axes ("generic neighbor/type pressure + small residuals"), which the readout later separates by sign/magnitude at the graph level. GAT shows peaks that are present but typically smaller; attention can separate channels earlier, while GCN produces smoother, less volatile curves.
The WNO/SI peak typically coincides with saturation of the number of active features/probes and the onset of a flat test loss. This aligns with two competing effects of more width: feature formation (easier to instantiate additional task-useful directions) versus packing efficiency (easier to spread existing directions). Early on, packing dominates; around $d\!\approx\!k$ capacity is spent on forming new directions (SI/WNO rise); once formation saturates, packing dominates again and SI/WNO drop. Centroids are tied directly to the BCE objective, so their feature formation saturates earlier, explaining the shorter first phase.
4.3 Topology-induced superposition
We train 3-layer GCN, GIN, and GATv2 models (16 channels per layer), apply a global $\mathsf{mean}$ pool followed by a ReLU, and fit a two-logit readout $(y_{A},y_{B})$. For each architecture we repeat training over 250 seeds. We study two feature families: (i) node-level probes $\mathsf{Inside}(C_{\ell})$ on the last pre-pooling layer, and (ii) graph-level probes $\mathsf{Has}(C_{\ell})$ on the post-pooling representation.
Figure 2: SI and WNO measured on GNNs trained on CONJUNCTION. Red highlighted data-points indicate models with perfect test accuracy.
Despite $d\!\gg\!k$, superposition remains large. With k=4k{=}4 motifs and width d=16d{=}16, independent axes would yield $\text{SI}\!\approx\!1$ and $\text{WNO}\!\approx\!0$. Instead, Fig. 2 shows clusters at $\text{SI}\!>\!1$ and $\text{WNO}\!\gg\!0$, i.e., above random overlap. In most runs, at least one motif pair has $|\cos|>0.9$ (Table 1), most often $(C_{3},C_{4})$ or $(C_{5},C_{6})$; in rare cases ($\sim$1% in GIN) all four are nearly collinear, consistent with an emergent lever of ’cycleness’. High accuracy can still be achieved, suggesting that the pooling / reading takes advantage of residual multicoordinate structure or magnitude even when node-level directions coalesce.
Cosine geometry reveals a cyclic structure. Figure 3 reports mean cosine similarity matrices. At the node level, the lengths of the nearby cycles align (for example, $C_{3}$ with $C_{4}$, $C_{5}$ with $C_{6}$) while distant pairs are weakly aligned or slightly antialigned. An MPNN layer $L$ implements a low-degree polynomial in the normalized adjacency, so detectors for $C_{3}\!\ldots C_{6}$ reuse monomials of closed walks up to length $\approx 2L{+}1$, naturally making adjacent lengths more similar. After pooling, geometry re-mixes toward the task: the post-pool ReLU and linear readout learn graph-level axes that are linear combinations of node-level detectors (e.g., in GIN with mean-pool, $C_{3}$ aligns with $C_{6}$ while $C_{4}$ opposes $C_{5}$, which helps reject single-motif false positives).
Figure 3: Mean cosine similarities for the node features (top) and graph features (bottom).
Node Graph Model Pool all one all one GCN mean 15.1% 87.8% 0.0% 44.9% GCN max 6.7% 78.8% 0.0% 20.2% GIN mean 1.1% 44.4% 0.0% 1.1% GIN max 0.4% 41.2% 0.0% 9.2% GAT mean 2.4% 64.1% 0.0% 10.0% GAT max 0.7% 63.0% 0.0% 12.2%
Table 1: Percentage of instances above 0.9 absolute cosine similarity.
Max pooling reduces lever sharing. Since a linear probe’s sign can flip with an equivalent boundary, we track $|\cos|$ to detect shared levers. Switching from mean to max pooling reduces node-level alignment across all architectures: both the fraction of runs with any $|\cos|>0.9$ pair and with all pairs >0.9>0.9 drop (Table 1). Intuitively, a per-channel max cannot encode a conjunction on a single channel – one large activation would mask the other—so the model is pressured to place motifs on distinct channels, lowering $|\cos|$. The ordering in Table 1 (GCN >> GAT >> GIN in lever sharing) aligns with each layer’s mixing bias. GCN’s degree normalized sum acts as a low-pass smoother, concentrating signals into a few graph-harmonic axes and driving different motifs to share directions. GAT’s learned attention reduces indiscriminate mixing and yields moderate de-alignment, but a single head offers limited selectivity. GIN’s unnormalized sum plus post aggregation MLP provides the most expressive channel rotation and gating, producing more specialized directions and the lowest node and graph level alignment.
Discussion. Even with generous width, semantically different topology features occupy similar node-space directions. Max pooling alleviates but does not eliminate this, probably reflecting message-passing inductive biases. Note that many high-accuracy models sit at $\text{SI}>1$ and $\text{WNO}\gg 0$ (Fig. 2), indicating that accurate solutions often rely on node-level lever sharing with disentanglement deferred to the graph-level readout – superposition here is not merely a symptom of under-capacity.
5 Pooling-Induced Axis Alignment of Graph Representations
Metric and preprocessing. We measure axis preference with the Alignment Index (AI) computed on any set of feature directions $\{c_{\ell}\in\mathbb{R}^{d}\}_{\ell=1}^{k}$ (node-level probe normals or graph-level class features):
$\mathrm{AI}\;=\;\frac{1}{k}\sum_{\ell=1}^{k}\frac{\max_{j}|(c_{\ell})_{j}|}{\|c_{\ell}\|_{2}}.$
Thus, $\mathrm{AI}\approx 1/\sqrt{d}$ for random orientations and $\mathrm{AI}\to 1$ when each feature is concentrated on a coordinate axis. We always use row-unit directions; for centroids we COM-center across classes before normalizing (to remove the shared offset), while probes use the raw fitted normals.
Global and local pooling conditions. For the graph readout we use the signed power-mean family that interpolates mean ($p=1$) and max ($p\to\infty$):
$y_{d}\;=\;\operatorname{sgn}(s_{d})\,N^{-1/p}\,|s_{d}|^{1/p},\qquad s_{d}=\sum_{i=1}^{N}\operatorname{sgn}(x_{id})\,|x_{id}|^{p}.$
To isolate local effects, we also replace the last message-passing aggregator by the same power-mean operator while holding the global readout at mean pooling.
Alignment vs. pooling and regime. Across both a bottleneck ($d=10$) and a wide ($d=22$) regime, AI increases as p>1p\!>\!1 (Fig. 4). With global pooling the effect is clearest at the graph level (the nonlinearity acts after node mixing). With a local power-mean in the last layer, AI increases at both node and graph levels because the winner-take-most bias precedes the final ReLU and readout. Occasionally AI dips for p>2p\!>\!2 in the bottleneck regime; this coincides with less stable training and noisier features. The trends persist when conditioning on high-accuracy runs. GATv2 tends to exhibit higher node-level AI even for modest $p_{\text{local}}$ (attention already induces axis-selective flows), whereas GCN’s degree-normalized linear aggregation produces smoother trends.
From alignment to constrained superposition. Under mean pooling, strong bottlenecks yield large $\mathrm{SI}$ due to shared directions, especially at the node level where separation is not directly optimized by the loss. Increasing $p_{\text{local}}$ reliably reduces $\mathrm{SI}_{\text{node}}$ (Fig. 4), effectively capping features-per-axis. Raising $p_{\text{global}}$ also lowers $\mathrm{SI}_{\text{node}}$ in the wide regime (consistent with §4.3); in the bottleneck regime the dimensional constraint dominates and this effect largely vanishes. $\mathrm{SI}_{\text{graph}}$ stays near-flat: in the wide regime it is already close to optimal under mean pooling (§4.2), and in the bottleneck regime additional alignment cannot overcome the shortage of dimensions.
(a) Wide; global
(b) Bottleneck; global
(c) Wide; local
(d) Bottleneck; local
Figure 4: Alignment and superposition indices versus pooling parameter $p$ across architectures.
Noise as the driver of axis alignment. Two complementary arguments explain why increasing $p>1$ breaks rotational symmetry in favor of axis alignment. (i) Learning-dynamics view. Backpropagating through power-mean pooling yields
$\frac{\partial h_{d}}{\partial x_{id}}\;=\;\frac{1}{N^{1/p}\,|s_{d}|^{1/p-1}}\operatorname{sgn}(s_{d})\,|x_{id}|^{p-1}\operatorname{sgn}(x_{id}),$
so large-magnitude coordinates get disproportionally large updates. With $p=1$ the Jacobian is rotation-invariant; with $p>1$ the extra factor $|x_{id}|^{p-1}$ biases gradient flow toward coordinates with larger magnitude, steadily aligning features to axes. In the absence of noise all orientations become equivalent once the model has recognized the features; in realistic graphs ubiquitous irrelevant features inject noise and trigger the symmetry breaking. (ii) Geometric view. Under equal-energy corruption, axis-aligned vectors lose less information than randomly oriented ones; in high dimensions the components of a random unit vector scale like $1/\sqrt{d}$ and are easily swamped by noise spikes. Figure 5 (from our diagnostic) visualizes this effect under max pooling. Empirically, purely noisy inputs confirm the variance story: max pooling inflates the standard deviation of pooled embeddings by an order of magnitude relative to mean pooling.
Practical takeaways. For more axis-aligned graph features, prefer global pooling with $p_{\text{global}}>1$. To make node channels less polysemantic, consider a power-mean local aggregator in the final layer.
Figure 5: Axes-aligned embedding vectors lose less information than arbitrary-angled vectors under max pooling when both are corrupted by equal-energy noise.
6 Rank collapse and metastable superposition
Our geometry metrics (SI/WNO) quantify how features pack inside the span of the learned representations. They do not say how many dimensions that span actually occupies. Let $H\!\in\!\mathbb{R}^{N\times d}$ stack pooled graph embeddings (rows are graphs) and let $C\!\in\!\mathbb{R}^{k_{a}\times d}$ stack the $k_{a}$ active class centroids. Since centroids are linear averages of rows of $H$, $C=A^{\top}HR$ for suitable averaging/projection matrices, hence
$\mathrm{EffRank}(C)\;\leq\;\mathrm{rank}(C)\;\leq\;\mathrm{rank}(H).$
If the span of $H$ is low-dimensional, superposition is inevitable: by pigeonhole, $k_{a}>\mathrm{rank}(H)$ forces multiple features to share axes. We formalize this with a numerical rank
$r_{\tau}(H)\;=\;\#\{i:\sigma_{i}(H)/\sigma_{1}(H)\geq\tau\}\!,$
and call a run collapsed when $r_{\tau}(H)<k_{a}$. [^3] We use $\tau\!=\!10^{-4}$; results are qualitatively unchanged for $\tau\in[10^{-5},10^{-3}]$. An energy criterion $r_{\eta}(H)=\min\{r:\sum_{i\leq r}\sigma_{i}^{2}\geq(1-\eta)\sum_{i}\sigma_{i}^{2}\}$ gives the same conclusions. This is the regime where high SI/WNO are not just a matter of suboptimal packing but a hard capacity constraint from a thin span.
Two training modes.
On PAIRWISE we observe two characteristic dynamics (Fig. 6). In some seeds $r_{\tau}(H)$ rises to the layer width (d=16d\!=\!16), with $\mathrm{EffRank}(H)$ tracking from below; SI/WNO on centroids drop accordingly. In other seeds $r_{\tau}(H)$ remains strictly below $d$, sometimes far below $k_{a}$: SI/WNO stay high and the smallest singular values show brief activation attempts (spikes) that recede, indicating the optimizer returns to a low-rank basin. Interestingly, across models that exhibit overfitting (such that the right model in Fig. 6) we observe sharp multi-phase transitions in SI and WNO throughout training. A possible (non-rigorous) explanation is provided in Appendix D.
Figure 6: Examples showing the evolution of singular values of a GCN model (left) and GIN model (right).
Why does the low-rank basin persist? Two effects make it metastable: (i) Global channel gating by ReLU. If a channel is negative across the dataset just before the last nonlinearity, ReLU zeros it everywhere, so the corresponding column of $H$ vanishes and algebraic rank drops exactly. Replacing the final ReLU with LeakyReLU substantially reduces such dead columns (App. E.6). (ii) Obtuseness pressure under BCE. Before a separating hyperplane forms, last-layer directions $\{\mathbf{v}_{i}\}$ tend to align with their class weights $\{\mathbf{w}_{i}\}$ and become mutually obtuse. Moving one feature into a fresh orthogonal dimension rotates it towards $90^{\circ}$ against the others and increases the BCE loss. After the hyperplane forms, the cross-class margins satisfy $\mathbf{w}_{j}^{\!\top}\mathbf{v}_{i}<0$ ($i\neq j$), and perturbing $\mathbf{v}_{k}$ into a new dimension makes those dots less negative, again increasing loss (Appendix E). Hence gradients oppose escapes until the accuracy gain from a new dimension outweighs the obtuseness penalty, explaining the singular jump events.
When is collapse likely? We see more collapse (i) with narrower last layers, (ii) in GIN (post-aggregation ReLUs) than GCN (no ReLU after the final aggregation), and (iii) as the number of task features grows. This agrees with the regime analysis in Appendix E: when $k_{a}\!\leq\!d+1$, mutually obtuse low-rank configurations exist and can act as local minima; as $k_{a}$ approaches or exceeds $2d$, such configurations become geometrically impossible and $r_{\tau}(H)$ tends to reach $d$.
Connection to oversmoothing. Classic oversmoothing is within-graph: node states become indistinguishable with depth. Here we see a cross-graph analogue: pooled graph embeddings concentrate in a thin global subspace. Accurate models can then separate classes inside that span by sign/magnitude, which explains why SI/WNO can be high at the node level yet acceptable in the graph space.
Discussion. The width-sweeps in §4.2 showed SI/WNO can be high even when $d\!\geq\!k$; here we identify a complementary mechanism: training can settle into thin-span solutions, so overlap is forced (because $k_{a}>r_{\tau}(H)$) rather than just suboptimal packing within a rich span. This also clarifies why node-level superposition can persist while graph-level centroids look well-packed: the readout operates inside a low-dimensional subspace, separating classes by sign/magnitude without increasing the span of $H$.
7 Conclusion
In this work, we provided, to the best of our knowledge, the first systematic study of superposition in GNNs. Concretely, we introduced a representation-centric framework that quantifies feature sharing in GNNs via class-conditional centroids and linear-probe directions. Across architectures, width induces a three-phase trajectory in overlap; topology shapes node-level geometry that pooling re-mixes toward the task; and sharper pooling drives axis alignment and reduces lever sharing. We also observed metastable low-rank graph embeddings in shallow models. Practically, modest increases in width, LeakyReLU in the final MLP, and sharper but stable pooling improve interpretability without sacrificing accuracy. A key next step is to link superposition dynamics to over-smoothing and over-squashing, integrating our geometric view with spectral, dynamical, and topological analyses.
Appendix A Generalized Mean Pooling
Let $N$ be the number of nodes, $x_{id}\in\mathbb{R}$ the $d$-th dimension of node $i$, and $p\in\mathbb{R}$ a pooling parameter. Introduce a small stabiliser $\varepsilon>0$ and define
$g_{p,\varepsilon}(x)\;=\;\operatorname{sgn}(x)\,\bigl{(}|x|+\varepsilon\bigr{)}^{p}.$
1. Element-wise transformation: For every node $i$ and feature dimension $d$,
$\tilde{x}_{id}=g_{p,\varepsilon}(x_{id})=\operatorname{sgn}(x_{id})\bigl{(}|x_{id}|+\varepsilon\bigr{)}^{p}.$
2. Pooling: Sum the transformed features and take their mean,
$s_{d}\;=\;\sum_{i=1}^{N}\tilde{x}_{id},\qquad\bar{s}_{d}\;=\;\frac{1}{N}\,s_{d}.$
3. Inverse transformation: Apply the signed $p$-th root to obtain the pooled feature
$h_{d}\;=\;\operatorname{sgn}(\bar{s}_{d})\,\bigl{(}|\bar{s}_{d}|+\varepsilon\bigr{)}^{1/p}.$
The resulting graph-level representation is the vector
$\mathbf{h}=\bigl{(}h_{1},h_{2},\dots,h_{D}\bigr{)}.$
Remark. Setting $p=1$ recovers ordinary mean pooling (modulo the stabilisation term $\varepsilon$), while $p\to\infty$ recovers max pooling.
Appendix B Symmetry Breaking
From Appendix A we note that the pooled elements are of the form:
$h_{d}\;=\;\operatorname{sgn}(\bar{s}_{d})\,\bigl{(}|\bar{s}_{d}|+\varepsilon\bigr{)}^{1/p}.$
Applying the chain rule gives:
$\frac{\partial h_{d}}{\partial x_{id}}=\frac{1}{N^{1/p}}\lvert s_{d}\rvert^{\frac{1}{p}-1}\operatorname{sgn}(s_{d})\times\lvert x_{id}\rvert^{p-1}\operatorname{sgn}(x_{id}).$
Appendix C SVD-Analysis
Let $H\in\mathbb{R}^{n\times d}$ be the matrix whose columns collect the averaged $d$-dimensional pooled embeddings of single-active-feature graphs, where $n$ is the number of features. Performing a singular–value decomposition,
$H\;=\;U\,\Sigma\,V^{\top},$
we obtain
  • $U\in\mathbb{R}^{n\times n}$: an orthogonal matrix whose columns $u_{1},\dots,u_{n}$ are the left singular vectors $\bigl{(}U^{\top}U=I_{n}\bigr{)}$;
  • $\Sigma\in\mathbb{R}^{n\times d}$: a (generally rectangular) diagonal matrix
    $\Sigma=\operatorname{diag}(\sigma_{1},\dots,\sigma_{r},0,\dots,0),\qquad\sigma_{1}\geq\sigma_{2}\geq\dots\geq\sigma_{r}>0,$
    where $r=\operatorname{rank}(H)$ and the positive entries $\sigma_{i}$ are the singular values;
  • $V\in\mathbb{R}^{d\times d}$: an orthogonal matrix whose columns $v_{1},\dots,v_{d}$ are the right singular vectors $\bigl{(}V^{\top}V=I_{d}\bigr{)}$.
Interpretation for pooled embeddings.
In our context, the directions $v_{i}$ in ascending order capture the most dominant linear subspaces along which the pooled embeddings vary when a single graph feature is activated. Hence, if any $\sigma_{i}=0$, the embeddings lie in a lower dimensional subspace.
Appendix D Sharp SI transitions during transition
A training-time phase transition.
On PAIRWISE we observe two training modes: (i) full-rank runs where $r_{\tau}(H)$ rises to $d$ and $\mathrm{EffRank}(H)$ tracks it; (ii) low-rank runs where $r_{\tau}(H)$ remains <d<d for most of training. Across seeds, SI/WNO of the centroids exhibit a reproducible four-stage pattern:
  1. 1.
[…]
  2. 2.
    Compression (sharp transition). A narrow span forms: smallest singular values of $H$ drop or remain suppressed, PC1 energy rises, and $\mathrm{EffRank}(C)$ contracts $\Rightarrow$ a sharp SI spike. This point almost always coincides with the onset of overfitting (test loss starts to rise).
  3. 3.
[…]
  4. 4.
    Convergence (flat). Geometry stabilizes; train loss drifts down, test loss drifts up.
The sharpness of Stage 2 reflects a discrete change in $r_{\tau}(H)$ (entire channels becoming globally inactive or reactivated), whereas WNO is computed after PC1 removal and need not jump at the same epoch.
Appendix E Geometric Regimes for Anti-Aligned Embedding Directions
E.1 Problem Formulation
We first consider the case where the feature number $n$ is small and no hyperplane has formed: Let $\{\,\mathbf{v}_{i}\in\mathbb{R}^{d}\mid\|\mathbf{v}_{i}\|_{2}=1,\;i=1,\dots,n\}$ be the (column-normalized) embedding directions learned by the network’s last hidden layer, and let $\{\mathbf{w}_{i}\}_{i=1}^{n}$ denote the corresponding unit-length rows of the classifier weight matrix. The anti-alignment hypothesis we explored in Section 6 states that after training
$\mathbf{v}_{i}\;\approx\;\mathbf{w}_{i}\qquad\text{and}\qquad\mathbf{v}_{i}^{\!\top}\mathbf{v}_{j}\;\leq\;0\quad(i\neq j),$
i.e. every embedding is (almost) identical to its class weight, and pairwise inner products are non-positive if possible. In other words, the network attempts to realise a set of $n$ mutually obtuse vectors in $\mathbb{R}^{d}$. The feasibility of (A.1) depends on the relation between $n$ and $d$, which can be separated into four distinct regimes.
E.2 Regime Analysis
  1. 1.
    Over-complete case, $n>2d$. Via a simple counting argument, at least two vectors must lie in the same closed half-space; hence their inner product is >0>0, so one pair is necessarily acute.
  2. 2.
    Intermediate regime, $d+1<n<2d$. There are at most $n=d+1$ mutually obtuse vectors. In this regime some vectors are mutually orthogonal and some are mutually obtuse.
  3. 3.
    Simplex threshold, $n=d+1$. The largest set of vectors that can satisfy $\mathbf{v}_{i}^{\!\top}\mathbf{v}_{j}<0$ for all $i\neq j$ is the vertex set of a centred regular $d$-simplex. This is the first value of $n$ that permits pairwise obtuse directions.
  4. 4.
    Under-complete case, $n\leq d$. One can choose any $n$ orthogonal unit vectors (or the $n$ vertices of a degenerate $n$-simplex in a n1n-1 dimensional subspace).
E.3 Implication for Rank Collapse
Assume training has converged to a low-rank subspace $\mathcal{S}\subset\mathbb{R}^{d}$ with $\dim(\mathcal{S})=r<d$ and all $\mathbf{v}_{i}\in\mathcal{S}$. Activating a new latent dimension $\mathbf{e}_{\perp}\perp\mathcal{S}$ perturbs one vector as $\mathbf{v}_{k}\mapsto\tilde{\mathbf{v}}_{k}=\sqrt{1-\varepsilon^{2}}\,\mathbf{v}_{k}+\varepsilon\mathbf{e}_{\perp}$. Then, for $i\neq k$,
$\tilde{\mathbf{v}}_{k}^{\!\top}\mathbf{v}_{i}=\sqrt{1-\varepsilon^{2}}\,\mathbf{v}_{k}^{\!\top}\mathbf{v}_{i}\;\geq\;\mathbf{v}_{k}^{\!\top}\mathbf{v}_{i},$
because $\mathbf{v}_{k}^{\!\top}\mathbf{v}_{i}\leq 0$ in the metastable state. Hence the angle decreases [^4] Equality holds only when $\mathbf{v}_{k}^{\!\top}\mathbf{v}_{i}=0$, keeping the angle unchanged at $90^{\circ}$. towards $90^{\circ}$, which reduces pairwise obtuseness. As discussed in Section 6 this is is penalised by the BCE loss function. During back propagation gradients oppose the escape from the low-rank basin, effectively creating a metastable minima. If $\varepsilon$ grows large enough performance gains may begin to outweigh the angle penalty, and the rank grows.
E.4 Rank Collapse with Hyperplane
So far we have assumed that
$\mathbf{v}_{i}\;\approx\;\mathbf{w}_{i},$
yet when a separating hyperplane emerges, the learned embeddings $\{\mathbf{v}_{i}\}$ and the classifier weights $\{\mathbf{w}_{i}\}$ are no longer almost identical. Empirically, $\mathbf{v}_{i}$ are often acute rather than mutually obtuse, but the cross-class dot products stay strictly negative:
$\mathbf{w}_{j}^{\!\top}\mathbf{v}_{i}\;<\;0\qquad(i\neq j).$
Perturbation into a fresh dimension.
Using the same notation as before, we perturb one embedding as
$\tilde{\mathbf{v}}_{k}\;=\;\sqrt{1-\varepsilon^{2}}\,\mathbf{v}_{k}+\varepsilon\,\mathbf{e}_{\perp}.$
For any $j\neq k$
$\mathbf{w}_{j}^{\!\top}\tilde{\mathbf{v}}_{k}\;=\;\sqrt{1-\varepsilon^{2}}\,\mathbf{w}_{j}^{\!\top}\mathbf{v}_{k}+\varepsilon\,\mathbf{w}_{j}^{\!\top}\mathbf{e}_{\perp}\;>\;\mathbf{w}_{j}^{\!\top}\mathbf{v}_{k},$
since $\mathbf{w}_{j}^{\!\top}\mathbf{v}_{k}<0$ by (A.4) while $\mathbf{w}_{j}^{\!\top}\mathbf{e}_{\perp}=0$. Thus the cross-class dot product becomes less negative, which increases BCE loss. In short, even after the hyperplane forms, the network remains trapped in a low-rank metastable basin for the same reason as before: cross-class obtuseness ($\mathbf{w}_{j}^{\!\top}\mathbf{v}_{i}<0$) resists escapes that would reduce the margin.
E.5 Singular-Value Thresholds
Here we make a few general remarks about the threshold for singular values, which signals the end of the metastable regime.
  1. 1.
    We expect from our theoretical description in Appendix E.3, that the threshold value depends on both the number of active features and the rank of the associated representation. This is because the more obtuse active features are, the stronger the effect that creates the energy barrier.
  2. 2.
    We empirically find the threshold to be approximately $10^{-4}$, based on a sample size of 50.
  3. 3.
    The threshold characteristics change noticeably with the active feature number and the rank of the associated representation.
E.6 Leaky ReLU Rank
Figure 7 illustrates the evolution of the rank of pooled embeddings across 15 training runs of a GIN model on the pairwise dataset. The left graph presents results for a standard GIN, while the right graph shows outcomes when the MLP layers in the GIN architecture use leaky ReLU as the activation function.
Figure 7: GIN model training runs with input dimension 12 and hidden dimensions [12, 6].